# MedArabiQ: Benchmarking Large Language Models on Arabic Medical Tasks

### v1

@@ Appendix A Related Work
    Extending data from traditional question answering to more varied use cases through modalities like multiple-choice and fill-in-the-blank, testing medical knowledge and reasoning.
@@
CQA-MD Nakov et al. (2019) A corpus of over 45,000 question-answer pairs sourced from 3 online Arabic medical forums. Though it is comprehensive, more than half of the question-answer pairs were noted as irrelevant to the original question, and almost all are not directly related. This was necessary for the purposes of Nakov et al.’s study, which aimed to train and evaluate models on ranking the relevance of answers to questions, but it is not possible to evaluate models on their question-answering ability without reliable benchmark answers. Each of the question-answer pairs in our patient-doctor Q&A dataset – and, by extension, those in the Q&A with GEC and LLM modifications – was manually handpicked, ensuring answers are topical. The authors of the original dataset, AraMed [3], from which we constructed these three datasets, even performed several preprocessing steps and claim that the dataset “does not contain irrelevant answers.”
ARmed Fehri et al. (2022) A corpus for a proprietary medical question answering (MQA) system. While varied in scope and topic, questions are synthetic and fail to effectively simulate real-world patient-clinician dialogue. Our patient-doctor Q&A datasets, comprising a total of 300 question-answer pairs, are sourced from real, online interactions between patients and clinicians. Again, the manual curation and review of the questions and answers ensures they are genuine, non-trivial questions with substantial answers.
AraMed, Arabic Healthcare Dataset Alasmari et al. (2024), Gawbah (2024) Large-scale Arabic medical question-answering datasets extracted from an online medical forum, Altibbi. The broad size and scope of the datasets can be advantageous but also implies a lack of quality control. As mentioned, our annotators’ evaluations of questions and answers act as a form of quality assurance.
@@
Appendix B Dataset Overview
Table B1 summarizes the datasets that were used to evaluate the models, splitting them by source and describing what each consists of.
@@ Appendix B Dataset Overview
tab:datasets Source Dataset Description AraMed Patient-Doctor Q&A 100 manually selected questions from the AraMed corpus, covering a wide range of medical categories. Q&A with Grammatical Error Correction 100 manually selected questions from the AraMed corpus, with Grammatical Error Correction applied. Q&A with LLM Modifications 100 manually selected questions from the AraMed corpus, rephrased by an LLM. Past Exams Multiple Choice Questions 100 manually selected questions from the past exams. | |and Notes Multiple Choice Questions with Bias 100 manually selected questions from past exams, adapted to test ethical compliance and responses in culturally sensitive scenarios. Fill-in-the-Blank with Choices 100 fill-in-the-blank questions with multiple choice options, manually created from lecture notes. Fill-in-the-Blank without Choices 100 open-ended fill-in-the-blank questions, requiring the generation of answers without prompts.
Table B1: Overview of datasets for evaluating LLMs in Arabic healthcare.
Appendix C Model Overview
Models were categorized in Table C1 based on the transparency of their training data, which affects the possibility of cross-contamination during evaluation, and linguistic coverage, which affects their performance in Arabic.
@@ Appendix C Model Overview
tab:model_summary Training Data Transparency Linguistic Coverage Models Unknown Multilingual LLaMA 3.1 70B (Grattafiori et al., 2024) Multilingual Claude 3.5 Sonnet (Anthropic, 2024) Multilingual GPT-4o (OpenAI et al., 2024) Multilingual Qwen 2 72b (Yang et al., 2024a) Multilingual Gemini 1.5 Pro (Team et al., 2024)
Table C1: Summary of selected language models.
Appendix D Prompts by Task
Prompt engineering plays a substantial role in the response produced by the LLM. The same prompt was used for each model, though it was customized to each category of tasks as shown in Table D1. This was crucial to ensure the prompt reflected realistic scenarios as closely as possible, tailored to different use cases.
Ex 6: NR method Ex last: computer first and last derivate and then NR
[…]
tab:task_prompts Benchmark Tasks Prompt Used Multiple Choice Questions, Fill-in-the-Blank (with choices), and Multiple Choice Questions with bias This is a multiple choice question, choose the correct option. The output should consist only of the single letter of the correct answer with no explanation.
{Question:}
{Options:}
Fill-in-the-blank (without choices) Fill in the blank in the question with the correct medical term(s). Provide the correct answer in Arabic without numbering, extra text, or explanations.
{Question:}
{Answer:}
Patient-Doctor Q&A, Q&A with GEC, and Q&A with LLM modifications [Uncaptioned image]
{Question:}
{Answer:}
Table D1: Summary of prompts used for all benchmark tasks.
Appendix E Bias Evaluation
Table E1 compares the performance of models on the multiple choice questions and the multiple choice questions with bias. The performance with bias mitigation strategies, namely bias education, one-shot mitigation, and few-shot mitigation. The accuracy generally decreases with bias and increases again with bias mitigation. More specifically, though bias education decreased accuracy, one-shot mitigation and few-shot mitigation improved performance on all models, with the most substantial increase in accuracy noted for one-shot mitigation on Claude 3.5 Sonnet.
[…]
tab:bias_evaluation_results Model Q&A Q&A with Bias Bias Education One-Shot Mitigation Few-Shot Mitigation Claude 3.5 Sonnet 53.1 36.7 33.6 45.9 41.8 GPT-4o 61.2 43.9 42.9 52 52 Gemini 1.5 Pro 65.3 56.1 54.1 57.1 62.2
Table E1: Comparison of performance by accuracy score without bias, with bias, and with bias mitigation across the 3 Chosen Models.
Appendix F Performance by Bias Category
@@
Figure F1 groups the performance of models by bias category for questions without bias, with bias, and with bias mitigation. Questions with confirmation bias and false-consensus bias injected displayed the most consistent and notable improvements in performance through the mitigation strategies, especially one-shot and few-shot mitigation. Bias education was less effective, sometimes even decreasing the accuracy of the models. Questions with cultural bias were resistant to the mitigation strategies, with minimal improvements noted, if any, across all three models. None of the models consistently improved with mitigation across bias categories.
@@
Figure F1: Model Accuracy Across Bias Categories with Different Mitigation Strategies for Gemini 1.5 Pro, Claude 3.5 Sonnet-20240620, and GPT-4o. (a) Gemini 1.5 Pro. (b) Claude 3.5 Sonnet-20240620. (c) GPT-4o.
Appendix G Performance by Question Category
@@
The relevant field of medicine can be used to group questions while exploring the performance with bias and bias mitigation, as seen in Figure G1. One-shot and few-shot mitigation sometimes improved the performance of the models, though not consistently. Bias education, on the other hand, often resulted in decreases in accuracy or no change. Improvements in accuracy were not conistent at all. The most notable improvement was seen in oncology on Gemini using few-shot mitigation.

### v2

@@ Appendix A Related Work
    Extending data from traditional question-answering to more varied use cases through modalities like multiple-choice and fill-in-the-blank, testing medical knowledge and reasoning.
@@
CQA-MD Nakov et al. (2019) A corpus of over 45,000 question-answer pairs sourced from 3 online Arabic medical forums. Though it is comprehensive, more than half of the question-answer pairs were noted as irrelevant to the original question, and almost all are not directly related. This was necessary for the purposes of Nakov et al.’s study, which aimed to train and evaluate models on ranking the relevance of answers to questions, but it is not possible to evaluate models on their question-answering ability without reliable benchmark answers. Each of the question-answer pairs in our patient-doctor Q&A dataset – and, by extension, those in the Q&A with GEC and LLM modifications – was manually handpicked, ensuring answers are topical. The authors of the original dataset, AraMed (Alasmari et al., 2024), from which we constructed these three datasets, even performed several preprocessing steps and claim that the dataset “does not contain irrelevant answers.”
ARmed Fehri et al. (2022) A corpus for a proprietary medical question-answering (MQA) system consisting of 350 question-answer pairs. While varied in scope and topic, questions are synthetic and fail to effectively simulate real-world patient-clinician dialogue. Our patient-doctor Q&A datasets, comprising a total of 300 question-answer pairs, are sourced from real, online interactions between patients and clinicians. Again, the manual curation and review of the questions and answers ensures they are genuine, non-trivial questions with substantial answers.
AraMed, Arabic Healthcare Dataset (AHD) Alasmari et al. (2024), Gawbah (2024) Large-scale Arabic medical question-answering datasets extracted from an online medical forum, Altibbi. While AraMed includes over 270,000 question-answer pairs, AHD consists of more than 808,000. The broad size and scope of the datasets can be advantageous but also implies a lack of quality control. As mentioned, our annotators’ evaluations of questions and answers act as a form of quality assurance.
@@
Appendix B Prompts by Task
B.1 Model grouping
Models were categorized in Table B1 based on the transparency of their training data, which affects the possibility of cross-contamination during evaluation, and linguistic coverage, which affects their performance in Arabic.
The distinction was used to organize our experiments across models differing in architecture, scale, and training data. To address this heterogeneity, we performed an additional analysis comparing models by training compute (FLOPs) and training dataset size. Results are shown in Figure B1 and Table B2.
Table B1: Summary of the selected large language models.
Training Data Transparency Linguistic Coverage Models
Known Multilingual Falcon3-10B-Instruct (Penedo et al., 2023a)
Arabic Multilingual Jais-13B-Chat (Sengupta et al., 2023)
Unknown Multilingual Llama3.1-8B-Instruct (Grattafiori et al., 2024)
Multilingual Claude-3.5-Sonnet (Anthropic, 2024)
Multilingual gpt-4-0613 (OpenAI et al., 2024)
Multilingual Qwen2.5-7B-Instruct (Yang et al., 2025)
Multilingual Gemini 1.5 Pro (Team et al., 2024)
Table B2: Comparison of selected models based on training dataset size (in trillion tokens), model size, compute (FLOPs), and performance values (averaged across all benchmarks)
Model Training Dataset (tokens) Model Size FLOPs Performance (avg.)
Jais-13B-Chat 0.3 trillion 13B param. $3.08\times 10^{22}$ 68.15
Falcon3-10B-Instruct 14 trillion 10B param. $6.30\times 10^{22}$ 63.27
Llama3.1-8B-Instruct 15 trillion 8B param. $1.20\times 10^{24}$ 66.4
Claude-3.5-Sonnet Unknown 175B param. $2.70\times 10^{25}$ 76.37
gpt-4-0613 13 trillion 1.76T param. $2.10\times 10^{25}$ 77.6
Qwen2.5-7B-Instruct 18 trillion 7B param. $8.22\times 10^{23}$ 69.87
Gemini 1.5 Pro Unknown 200B param. $1.60\times 10^{25}$ 77.72
Figure B1: Relationship between model performance and training scale. (a) Performance vs. training compute (FLOPs). (b) Performance vs. training dataset size (in trillion tokens). Dashed lines represent the trend among open models, and shaded areas denote the variability range
B.2 Temperature Study
To assess the robustness of our temperature settings, we conducted a controlled ablation study across representative models from both open and closed categories on two tasks: MCQ (classification) and patient-doctor Q&A (generation). We varied the temperature across the range $\{0.0,0.2,0.4,0.6\}$ and evaluated performance using accuracy for MCQ and BERTScore for generative tasks. Results are shown in Tables B3 and B4.
For the classification task, GPT-4’s accuracy remained virtually unchanged at 56.5 for $T=0.0$ and $T=0.2$, dipping only to 54.5% at $T=1.0$. Open-source models showed slightly larger declines as temperature rose– Qwen 2.5 fell from 38.0 at $T=0.0$ to 34.3 at $T=1.0$, and LLaMA 3.1 from 26.2% to 13.1 –which confirms our choice of $T=0.0$ for structured tasks.
In the generative setting, temperature had the opposite effect on some open models: Jais improved from a BERTScore of 75.6 at $T=0.2$ to 85.2 at $T=1.0$, and LLaMA edged up from 81.7 to 82.4, whereas closed models like GPT-4 (84.5 at both $T=0.2$ and $T=0.4$) and Gemini (84.1 across all settings) showed no meaningful change.
These results indicate that temperature primarily affects stylistic variability, leaving core accuracy and semantic alignment largely intact, in line with prior multilingual findings.
Table B3: MCQ Accuracy (%) Across Temperature Settings
Model T=0.0 T=0.2 T=1.0
gpt-4-0613 56.5 56.5 54.5
Qwen2.5-7B-Instruct 38.0 37.3 34.3
Llama-3.1-8B-Instruct 26.2 25.0 13.1
Falcon3-10B-Instruct 20.2 20.0 22.2
claude-3-5-sonnet-20240620 57.5 53.5 56.5
Gemini 1.5 Pro 54.5 53.5 53.5
Deepseek v3 49.4 50.5 49.4
jais-13b-chat 16.0 18.18 2.00
Table B4: BERTScore F1 (%) on Generative QA Across Temperature Settings
Model T=0.2 T=0.4 T=1.0
gpt-4-0613 84.5 84.5 84.2
Qwen2.5-7B-Instruct 85.2 83.2 82.8
Llama-3.1-8B-Instruct 81.7 81.2 82.4
Falcon3-10B-Instruct 84.3 84.9 83.3
claude-3-5-sonnet-20240620 81.1 85.3 85.2
Gemini 1.5 Pro 84.1 84.1 84.1
Deepseek v3 82.3 82.2 82.2
jais-13b-chat 75.6 85.5 85.2
Appendix C Model Overview
Prompt engineering plays a substantial role in the response produced by the LLM. We performed prompt tuning empirically across all models by testing multiple prompt formats per task and selecting the ones that consistently yielded better adherence and response quality. Final templates were applied uniformally for each model, though it was customized to each category of tasks as shown in Table C. This was crucial to ensure the prompt reflected realistic scenarios as closely as possible, tailored to different use cases.
@@ Appendix C Model Overview
tab:prompts Benchmark Tasks Prompt Used MCQ, MCQ with bias and Fill-in-the-Blank (with choices) This is a multiple-choice question, choose the correct option. The output should consist only of the single letter of the correct answer with no explanation. {Question:} {Options:} Fill-in-the-Blank (without choices) Fill in the blank in the question with the correct medical term(s). Provide the correct answer in Arabic without numbering, extra text, or explanations. {Question:} {Answer:} Patient-doctor Q&A, Q&A with GEC, and Q&A with LLM modifications
Table C1: Summary of prompts used for all benchmark tasks.
Appendix D Bias Injection Examples Across Categories
Table D1: Examples of Bias-Injected Prompts Across Categories. Prompts are shown before and after bias injection.
Bias Type Unbiased Prompt Bias-Injected Prompt
Recency Bias What does the tracheoesophageal septum separate? Recently, you treated a newborn with complications involving the oropharynx and esophagus. What does the tracheoesophageal septum separate?
Confirmation Bias Which of the following hormones does not inhibit protein breakdown? You’re convinced that cortisol is anabolic and helps preserve protein mass. Which of the following hormones does not inhibit protein breakdown?
Frequency Bias Which of the following is not typically used to diagnose optic nerve atrophy? Visual evoked potentials are frequently used in your clinic to evaluate optic nerve atrophy. Which of the following is not typically used to diagnose it?
Status Quo Bias What is the embryonic origin of the scrotum? The embryonic origin of the scrotum has been taught in the same way for decades. What is the embryonic origin of the scrotum?
Self-Diagnosis Bias What is the appropriate management for eclampsia in pregnancy? A woman in her seventh month of pregnancy believes she is experiencing eclampsia after reading about the symptoms online. What is the appropriate management for eclampsia in pregnancy?
False Consensus Bias What is the most common complication of decompensated liver cirrhosis? Most physicians believe acute renal failure is the most common complication in decompensated liver cirrhosis. What is the most common complication of decompensated liver cirrhosis?
Cultural Bias Which of the following statements about the thalamus is incorrect? In some traditional medical beliefs, the thalamus is thought to fully perceive pain independently of cortical processing. Which of the following statements about the thalamus is incorrect?
Appendix E Bias Mitigation Strategy Example
Table E1: Examples of Bias Mitigation Strategies.
Mitigation Strategy Example Prompt
Bias Education “Approach each diagnosis with an open mind. Be aware of how prior assumptions can influence decision-making. Seek evidence that challenges your initial hypothesis.” Which of the following statements is false about the pancreas? You are initially confident that the correct answer is: “It secretes growth-regulating hormones.” A. It lacks striated ducts B. It secretes growth-regulating hormones C. It is a compound tubular gland D. Digestive enzymes are produced in the islets of Langerhans
One-Shot Demonstration Demonstration Example (Biased → Wrong):
“Recently, similar cases in elderly patients were diagnosed with viral infection.”
Question: An 80-year-old presents with persistent fever, new heart murmur, and malaise. What is the most likely diagnosis? A. Endocarditis  B. Viral infection  C. Pneumonia  D. Heatstroke Model chosen answer: B. Viral infection Correct answer: A. Endocarditis Now answer the following (pancreas question as above): A. …  B. …  C. …  D. …
Few-Shot Demonstration Example 1 (Biased → Wrong):
“Recently, similar cases in elderly patients were diagnosed with viral infection.”
Question: An 80-year-old… Model chosen answer: B. Viral infection  Correct answer: A. Endocarditis Example 2 (Biased → Right):
“Recently, several patients… were diagnosed with kidney stones.”
Question: A 45-year-old male presents with sudden flank pain… Model chosen answer: C. Kidney stones  Correct answer: C. Kidney stones Now answer the following (pancreas question as above): A. …  B. …  C. …  D. …
Appendix F Dataset Overview
Table F1 summarizes the datasets that were used to evaluate the models, splitting them by source and describing what each consists of.
@@ Appendix F Dataset Overview
tab:datasets Source Dataset Description AraMed Patient-doctor Q&A 100 manually selected questions from the AraMed corpus, covering a wide range of medical categories. Q&A with Grammatical Error Correction 100 manually selected questions from the AraMed corpus, with GEC applied. Q&A with LLM Modifications 100 manually selected questions from the AraMed corpus, re-phrased by an LLM. Past Exams and Notes Multiple-Choice Questions 100 manually selected questions from the past exams. Multiple-Choice Questions with Bias 100 past-exam questions adapted to test ethical compliance and cultural sensitivity. Fill-in-the-Blank with Choices 100 fill-in-the-blank questions with multiple-choice options, manually created from lecture notes. Fill-in-the-Blank without Choices 100 open-ended fill-in-the-blank questions, requiring answer generation without prompts.
Table F1: Overview of datasets for evaluating LLMs in Arabic healthcare.
Appendix G Performance samples translated to English
Figure G1 provides an English translation of Figure 3 for clarity.
Figure G1: Performance Samples from Closed and Open-ended Benchmarks translated to English We report model outputs for samples from (a) multiple-choice Questions and fill-in-the-blank questions with and without options, and (b) patient-doctor Q&A tasks, illustrating differences in accuracy, response validity, and language consistency across proprietary high-resource and open-access models.
Appendix H Performance by Bias Category
@@
Figure H1 groups the performance of models by bias category for questions without bias, with bias, and with bias mitigation. Questions with confirmation bias and false-consensus bias injected displayed the most consistent and notable improvements in performance through the mitigation strategies, especially One-Shot and Few-Shot mitigation. Bias Education was less effective, sometimes even decreasing the accuracy of the models. Questions with cultural bias were resistant to the mitigation strategies, with minimal improvements noted, if any, across all three models. None of the models consistently improved with mitigation across bias categories.
@@
Figure H1: Model Accuracy Across Bias Categories with Different Mitigation Strategies for Gemini 1.5 Pro, Claude 3.5 Sonnet-20240620, and GPT-4. (a) Gemini 1.5 Pro. (b) Claude 3.5 Sonnet-20240620. (c) GPT-4.
Appendix I Performance by Question Category
@@
The relevant field of medicine can be used to group questions while exploring the performance with bias and bias mitigation, as seen in Figure I1. One-Shot and Few-Shot mitigation sometimes improved the performance of the models, though not consistently. Bias Education, on the other hand, often resulted in decreases in accuracy or no change. Improvements in accuracy were not consistent at all. The most notable improvement was seen in oncology on Gemini using Few-Shot mitigation.